What is the mathematical symbol proof of

If A is a subset of B then P(B|A)=1 and P(A|B)=P(A)/P(B)
Answer #1

If A is a subset of B, then AnB = A.

Thus,

P(B|A) = P(AnB)/P(A) = P(A)/P(A) = 1

P(A|B) = P(AnB)/P(B) = P(A)/P(B)